OpenStudy (anonymous):

5 question quiz. Need help. Will fan and medal!!!

OpenStudy (anonymous):

1. Multiply and simplify. 6 1/4 and 4/5 A.1/5 B. 5 C. 1/4 D. 4

OpenStudy (anonymous):

2. multiply and simplify. 2/3* (1/3+ 3 1/4) A. 4 7/18 B. 4 7/9 C. 2 7/9 D. 2 7/18

OpenStudy (frozen15):

I can help on # 1 .. 6 1/4 = 25/4 * 4/5 = 100/20 = 5 Your answer is (b) I can help on # 2 .... 2/3 × (1/3 + 3 1/4) = 2/3 (1/3 + 13/4) = 2/3 (43/12) = 43/18 = 2 7/18 So your answer is (D) question #1 = B question #2 = D